Solve x y z and t in the matrix equation below [3x t+


1
Expert's answer
2021-05-20T17:17:36-0400

Given,

"\begin{bmatrix} 3x& y-x \\t+\dfrac{z}{2}& t-z\end{bmatrix} =\begin{bmatrix} 3& 1\\\dfrac{7}{2}& 3\end{bmatrix}"


On comparing we get-


"3x=3\Rightarrow x=1"


Also, "y-x=1\Rightarrow y=1+1\Rightarrow y=2"


Also,


"t+\dfrac{z}{2}=\dfrac{7}{2}~~~~~~~~~~~~~-(1)"


"t-z=3~~~~~~~~-(2)"


Solving eqn.(1) and (2) and we get-


"z=\dfrac{1}{3} , t=\dfrac{10}{3}"


Hence Value of "x=1,y=2,z=\dfrac{1}{3}" and "t=\dfrac{10}{3}"